### v2

@@ 4 Manifolds of absolutely continuous functions
is continuous and the point evaluation $(\varepsilon_{c})_{t}:C([a,b],N)\to N$, $\eta\mapsto\eta(t)$ is smooth for each $t\in[a,b]$ (see [1]). Since the inclusion map $J:AC_{L^{p}}([a,b],N)\to C([a,b],N)$ is smooth, we have $\varepsilon=\varepsilon_{c}\circ(J,\text{Id}_{\mathbb{R}})$ and $\varepsilon_{t}=(\varepsilon_{c})_{t}\circ J$ for each $t\in[a,b]$. ∎
@@
$F:[a,b]\times M\to N,\quad F(t,m)=f(\gamma(t),m)$
@@
For each $(t,m)\in[a,b]\times M$ we have
$F(t,m)=AC_{L^{p}}([a,b],f)\circ(\varepsilon(\gamma,\cdot)\times\text{id}_{M})(t,m)$
where $\varepsilon$ is the evaluation map as before. Therefore, the map $F$ is composition of continuous maps. ∎
@@
Let consider the local addition $\Sigma:\Omega\to N$ and $\theta_{N}:\Omega\to\Omega^{\prime}$ as in Definition 4.1. Let $(U,\varphi)$ be a chart in $q\in N$ such that $\{q\}\times U\subseteq\Omega^{\prime}$ and $(\mathcal{U}_{\zeta_{q}},\Psi_{\zeta_{q}}^{-1})$ be a chart in $\zeta_{q}\in AC_{L^{p}}([a,b],N)$.
@@
$\Theta_{\eta}:T_{\eta}AC_{L^{p}}([a,b],N)\to{\Gamma_{AC}(\eta)},\quad v\mapsto\Theta_{\eta}(v):=\Theta_{N}|_{T_{\eta}AC_{L^{p}}([a,b],N)}(v)$
@@
Let $\Sigma:\Omega\to N$ be a normalized local addition of $N$. Since ${\Gamma_{AC}(\eta)}$ is a vector space, we identify its tangent bundle with ${\Gamma_{AC}(\eta)}\times{\Gamma_{AC}(\eta)}$. Let $\Psi_{\eta}:\mathcal{V}_{\eta}\to\mathcal{U}_{\eta}$ be a chart on $\eta$ such that $\Psi_{\eta}(0)=\eta$, then
@@
is absolutely continuous with $\pi_{TN}\circ\Theta_{\eta}(v)=\eta$, making the map $\Theta_{\eta}$ an isomorphism of topological vector spaces. ∎
@@
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits local addition and $p\in[1,\infty]$. Since $TN$ admits local addition and the vector bundle $\pi_{TN}:TN\to N$ is smooth, the map
$AC([a,b],\pi_{TN}):AC_{L^{p}}([a,b],TN)\to AC_{L^{p}}([a,b],N),\quad\tau\mapsto\pi_{TN}\circ\tau$
is smooth. Moreover, if $\eta\in AC_{L^{p}}([a,b],N)$, then
$AC([a,b],\pi_{TN})^{-1}(\{\eta\})={\Gamma_{AC}(\eta)}.$
The following result follows the same steps as for the case of $C^{\ell}$-maps (with $\ell\geq 0$) from a compact manifold (possibly with rough boundary) to a smooth manifold which admits local addition [2, Theorem A.12].
@@
$AC([a,b],\pi_{TN}):AC_{L^{p}}([a,b],TN)\to AC_{L^{p}}([a,b],N),\quad\tau\mapsto\pi_{TN}\circ\tau$
is a smooth vector bundle with fiber ${\Gamma_{AC}(\eta)}$ over $\eta\in AC_{L^{p}}([a,b],N)$. Moreover, the map
@@
Let $M$ and $N$ be a smooth manifolds modeled on sequentially complete locally convex spaces which admits a local addition and $p\in[1,\infty]$. If $f:M\to N$ is a $C^{3}$-map, then the tangent map of
@@
By Proposition 4.8, since $f$ is $C^{3}$ we know that $AC_{L^{p}}([a,b],f)$ is $C^{1}$, thus $TAC_{L^{p}}([a,b],f)$ exists. Let consider the local addition $\Sigma_{M}:\Omega_{M}\to M$ and $\eta\in AC_{L^{p}}([a,b],M)$.
If $\Psi_{\eta}:\mathcal{V}_{\eta}\to\mathcal{U}_{\eta}$ is a chart on $\eta$ such that $\Psi_{\eta}(0)=\eta$. We consider the isomorphism of vector space
@@
$T\varepsilon_{t}^{N}\circ TAC_{L^{p}}([a,b],f)\circ T\Psi_{\eta}(0,\sigma)$ $=T\varepsilon_{t}\circ TAC_{L^{p}}([a,b],f)\circ T\Psi_{\eta}([s\mapsto s\sigma])$
@@
$=Tf\circ T_{0}\Sigma_{M}|_{T_{\eta(t)}N}([s\sigma(t)])$
@@
Let $p\in[1,\infty]$. If $G$ is a Lie group modeled in a sequentially locally convex space $E$, then we already know that the space $AC_{L^{p}}([a,b],G)$ is a Lie group with Lie algebra given by $AC_{L^{p}}([a,b],T_{e}G)$ (see [10, 24]). We will give an alternative proof of this.
Let $e\in G$ be the neutral element, let $L_{g}:G\to G$, $h\mapsto gh$ be the left translation by $g\in G$ and the action
$G\times TG\to TG,\quad(g,v_{h})\mapsto g.v_{h}:=TL_{g}(v_{h})\in T_{gh}G.$
@@
defines a local addition for $G$ (see e.g. [15]); hence $AC_{L^{p}}([a,b],G)$ is a smooth manifold with charts constructed with $(\Omega_{\varphi},\Sigma_{\varphi})$.
@@
This enable us to say that for the neutral element ${\zeta_{e}}\in AC_{L^{p}}([a,b],G)$ the chart is given by
@@
$TG\times G\to TG,\quad(v_{h},g)\mapsto v_{h}.g:=T\rho_{g}(v).$
And its restriction
$T_{e}G\times G\to TG,\quad(v,g)\mapsto v.g:=T\rho_{g}(v).$
Unlike on Lie groups, on half-Lie groups this actions may not be smooth. Hence we can not construct a local addition using a convenient chart on the identity (see Example 4.25).
@@
would imply the smoothness of the multiplication map on $G$. In fact, since the point evaluation map $\varepsilon_{a}:AC_{L^{p}}([a,b],G)\to G$, $\eta\mapsto\eta(a)$ and the map $\zeta:G\to AC_{L^{p}}([a,b],G)$, $g\mapsto[t\mapsto g]$, are smooth, the multiplication map on $G$ would be smooth as it coincides with the composition
@@
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space $E$ which admits local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. Let $P=\{t_{0},..,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ charts of $G$ such that verify the definition of absolutely continuous for $\eta$. For each $j\in\{1,...,n\}$ we denote
$\eta_{j}:=\varphi_{j}\circ\eta_{[t_{i-1},t_{i}]}\in AC_{L^{p}}([t_{i-1},t_{i}],E).$
Then $\eta_{j}^{\prime}\in L^{p}([t_{i-1},t_{i}],E)$. Let $\eta^{\prime}_{j}=[\gamma_{j}]$ with $\gamma_{j}\in\mathcal{L}^{p}([t_{i-1},t_{i}],E)$, we define $\gamma:[a,b]\to TG$ via
$\gamma(t):=T(\varphi_{j})^{-1}(\eta_{j}(t),\gamma_{j}(t)),\text{ if }t\in[t_{i-1},t_{i}[\text{ with }j\in\{1,...,n\},$
and $\gamma(b)=T(\varphi_{j})^{-1}(\eta_{j}(b),\gamma_{j}(b))$. Then $\gamma$ is measurable (see [10]) and we define
$\dot{\eta}:=\gamma.$
@@
${\dot{\eta}}_{\gamma}=\gamma.\eta_{\gamma},\quad\eta_{\gamma}(0)=e$
@@
$\operatorname{Evol}:L^{p}([0,1],T_{e}G)\to AC_{L^{p}}([a,b],G),\quad[\gamma]\mapsto\operatorname{Evol}(\gamma):=\eta_{\gamma}.$
@@
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space and $p\in[1,\infty]$. We say that $G$ is locally $L^{p}$-semiregular if there is exists an open 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that for each $[\gamma]\in B$ there exists a $AC_{L^{p}}$-Carathéodory solution $\eta_{\gamma}\in AC_{L^{p}}([0,1],G)$ of the equation
$\dot{\eta}_{\gamma}=\gamma.\eta_{\gamma},\quad\eta_{\gamma}(0)=e,$
@@
For our purpuse, we will use the subdivision property [24, Lemma 2.17].
@@
$\sup_{k\in\{1,...,n-1\}}q(\gamma_{n,k})\to 0,\quad\text{as }n\to\infty.$
@@
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. Then $G$ is locally $L^{p}$-semiregular if and only if $G$ is $L^{p}$-semiregular.
@@
Let $G$ be locally $L^{p}$-semiregular. Then there is exists and 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that $\operatorname{Evol}(\beta):=\eta_{\beta}\in AC_{L^{p}}([a,b],G)$ exists of each $[\beta]\in B$. without loss of generalization we assume that
@@
For $[\beta]\in W_{\gamma}$ and $j\in\{0,...,n-1\}$ we denote $\eta_{j}=\operatorname{Evol}(\beta_{n,j})$ and we define the function $\eta_{\beta}:[0,1]\to G$ via
@@
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. We say that $G$ is locally $L^{p}$-regular if $G$ is $L^{p}$-semiregular and there is exists an 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that its restricted evolution map $\operatorname{Evol}|_{B}$ is smooth.
The following Lemma is just an application of [23, Proposition 4.11] to our case.
@@
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. We consider the evolution map with continuous values
@@
Then, $\operatorname{Evol}_{C}$ is continuous if and only if there is exists an 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that its restricted evolution map $\operatorname{Evol}_{C}|_{B}$ is continuous.
@@
is continuous. Following Lemma 5.9, for each $[\beta]\in W_{\gamma}$, we have that the construction of the function $\operatorname{Evol}(\beta)$ implies that $Evol_{C}$ is just the product of composition of continuous maps. ∎
@@
Let $E_{1},E_{2}$ and $F$ be locally convex spaces, $1\leq p<\infty$ and $U\subseteq E_{2}$ an open subset. If $r\in\mathbb{N}\cup\{0,\infty\}$ and $f:E_{1}\times U\to F$ is a $C^{r}$-map such that for each $y\in U$ the map $f(\cdot,y):E_{1}\to F$ is linear, then the map
@@
Hence $\tilde{f}(\gamma_{0},\eta_{0})\in L^{p}([0,1],F)$. Since the space $C([0,1],E_{1})$ is dense in $L^{p}([0,1],E_{1})$, there exists a $\gamma_{c}\in C([0,1],E_{1})$ such that
$\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},p}\leq 2\varepsilon/5.$
@@
With this, for each $\eta\in W$ and $\gamma\in L^{p}([0,1],E_{1})$ with $\lVert\gamma-\gamma_{0}\lVert_{L^{p},\kappa}\leq 2\varepsilon/5$ we estimate
$\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}$ $\leq\lVert\tilde{f}(\gamma-\gamma_{c},\eta)\lVert_{L^{p},\beta}+\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}$
$\leq\lVert\tilde{f}(\gamma-\gamma_{c},\eta)\lVert_{L^{p},\beta}+\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\lVert_{L^{p},\beta}+\lVert\tilde{f}(\gamma_{c}-\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}$
$\leq\lVert\gamma-\gamma_{c}\lVert_{L^{p},\kappa}+\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\lVert_{L^{p},\beta}+\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},\kappa}.$
Since $\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},\kappa}\leq 2\varepsilon/10$ and $\lVert\gamma-\gamma_{0}\lVert_{L^{p},\kappa}\leq 2\varepsilon/10$ we have
$\lVert\gamma-\gamma_{c}\lVert_{L^{p},\kappa}\leq 2\varepsilon/5$
and
$\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}\leq 2\varepsilon/5+2\varepsilon/5+2\varepsilon/10=\varepsilon$
Thus $\tilde{f}$ is continuous. By linearity in the first variable the function $\overline{f}$ has a continuous differential $d_{1}\tilde{f}:L^{p}([0,1],E_{1})\times L^{p}([0,1],E_{1})\times C([0,1],E_{2})\to L^{p}([0,1],F)$ in the first variable. Let consider the map $f$ as a $C^{1}$-map. If $x\in E_{1}$, $y,y_{1}\in E_{2}$ and $t\in\mathbb{R}^{\times}$, then we have
@@
Hence the continuity of $\widetilde{d_{2}f}$ implies that the map $\overline{f}$ is $C^{1}$. Proceeding by induction, if $f$ is a $C^{k}$-map, since $\tilde{f}$ is linear in the first variable we have
@@
For the case $L_{rc}^{\infty}$ we recall [10, Proposition 2.3].
@@
Let $G$ be a right half-Lie group modeled in a Fréchet space $E$ which admits local addition and $1\leq p<\infty$. Let $G$ be $L^{p}$-semiregular with continuous evolution map
@@
is continuous.
If we replace $L^{p}$ with $L_{rc}^{\infty}$ the result remains valid.
@@
Then each $f_{i}$ is continuous and for each $x\in V_{i}$ the function $f_{i}(\cdot,x):T_{e}G\to E$, $v\mapsto f_{i}(v,x)$ is linear. Hence by Lemma 5.12 the map
@@
$F:\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i})\to\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],E),\quad\left([\beta_{i}],\eta_{i}\right)_{i=1}^{n}\mapsto\left(f_{i}\circ([\beta_{i}],\eta_{i}\right)_{i=1}^{n}.$
@@
Let denote the homeomorphism onto its image (see Proposition 3.9)

### v4

@@ 3 Manifolds of absolutely continuous functions
is continuous and the point evaluation $(\varepsilon_{c})_{t}:C([a,b],N)\to N$, $\eta\mapsto\eta(t)$ is smooth for each $t\in[a,b]$ (see [1]). Since the inclusion map $J:AC_{L^{p}}([a,b],N)\to C([a,b],N)$ is smooth, the assumptions follow from the observation that $\varepsilon=\varepsilon_{c}\circ(J\times\text{Id}_{[a,b]})$ and $\varepsilon_{t}=(\varepsilon_{c})_{t}\circ J$ for each $t\in[a,b]$. ∎
@@
$F:[a,b]\times M\to N,\quad(t,p)\mapsto f(\gamma(t),p)$
@@
It follows from the fact that $F=f\circ(\gamma\times\text{id}_{M})$. ∎
@@
Consider the local addition $\Sigma:\Omega\to N$ and $\theta_{N}:\Omega\to\Omega^{\prime}$ as in Definition 3.1. Let $(U,\varphi)$ be a chart around $q\in N$ such that $\{q\}\times U\subseteq\Omega^{\prime}$ and $(\mathcal{U}_{\zeta_{q}},\Psi_{\zeta_{q}}^{-1})$ be a chart in $\zeta_{q}\in AC_{L^{p}}([a,b],N)$.
@@
Following more general cases of manifolds of mappings, such as the case of $C^{\ell}$-maps (with $\ell\geq 0$) from a compact manifold (possibly with rough boundary) to a smooth manifold which admits local addition (see e.g. [3, 24]), we well study the tangent bundle of the manifold of absolute continuous functions.
@@
$\Theta_{\eta}:T_{\eta}AC_{L^{p}}([a,b],N)\to{\Gamma_{AC}(\eta)},\quad v\mapsto\Theta_{\eta}(v):=\Theta_{N}(v)$
@@
Let $\Sigma:\Omega\to N$ be a normalized local addition of $N$. Since ${\Gamma_{AC}(\eta)}$ is a vector space, we identify its tangent bundle with ${\Gamma_{AC}(\eta)}\times{\Gamma_{AC}(\eta)}$. Let $\Psi_{\eta}:\mathcal{V}_{\eta}\to\mathcal{U}_{\eta}$ be a chart around $\eta$ such that $\Psi_{\eta}(0)=\eta$, then
@@
is in $AC_{L^{p}}([a,b],TN)$ and verifies $\pi_{TN}\circ\Theta_{\eta}(v)=\eta$, making the map $\Theta_{\eta}$ an isomorphism of topological vector spaces. ∎
@@
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits local addition and $p\in[1,\infty]$. Since $TN$ admits local addition and the vector bundle $\pi_{TN}:TN\to N$ is smooth, by Proposition 3.8, the map
$AC_{L^{p}}([a,b],\pi_{TN}):AC_{L^{p}}([a,b],TN)\to AC_{L^{p}}([a,b],N),\quad\tau\mapsto\pi_{TN}\circ\tau$
is smooth. Moreover, for $\eta\in AC_{L^{p}}([a,b],N)$, we have
$AC_{L^{p}}([a,b],\pi_{TN})^{-1}(\{\eta\})={\Gamma_{AC}(\eta)}.$
Then $AC_{L^{p}}([a,b],\pi_{TN})$ defines a vector bundle on $AC_{L^{p}}([a,b],N)$.
By Proposition 3.8, we can apply [3, Theorem A.12] for our particular case:
@@
Let $M$ and $N$ be a smooth manifolds modeled on sequentially complete locally convex spaces which admits a local addition and $p\in[1,\infty]$. If $f:M\to N$ is a smooth map, then the tangent map of
@@
By Proposition 3.8, we know that $AC_{L^{p}}([a,b],f)$ is smooth, thus $TAC_{L^{p}}([a,b],f)$ makes sense. Let us consider the local addition $\Sigma_{M}:\Omega_{M}\to M$ and $\eta\in AC_{L^{p}}([a,b],M)$.
If $\Psi_{\eta}:\mathcal{V}_{\eta}\to\mathcal{U}_{\eta}$ is a chart around $\eta$ such that $\Psi_{\eta}(0)=\eta$. We consider the isomorphism of vector space
@@
$T\varepsilon_{t}^{N}\circ TAC_{L^{p}}([a,b],f)\circ T\Psi_{\eta}(0,\sigma)$ $=T\varepsilon_{t}^{N}\circ TAC_{L^{p}}([a,b],f)\circ T\Psi_{\eta}([s\mapsto s\sigma])$
@@
$=Tf\circ T_{0}\Sigma_{M}|_{T_{\eta(t)}M}([s\sigma(t)])$
@@
Let $p\in[1,\infty]$. If $G$ is a Lie group modeled in a sequentially locally convex space $E$, then we already know that the space $AC_{L^{p}}([a,b],G)$ is a Lie group (see [14, 33]). We will give an alternative proof of this.
Let $e\in G$ be the neutral element, let $L_{g}:G\to G$, $h\mapsto gh$ be the left translation by $g\in G$ and the left action
$G\times TG\to TG,\quad(g,v_{h})\mapsto g.v:=TL_{g}(v).$
@@
defines a local addition for $G$ (see e.g. [23]); hence $AC_{L^{p}}([a,b],G)$ is a smooth manifold with charts constructed with the local adition $(\Omega_{\varphi},\Sigma_{\varphi})$.
@@
This enables us to say that for the neutral element ${\zeta_{e}}\in AC_{L^{p}}([a,b],G)$ we have a chart given by
@@
$TG\times G\to TG,\quad(v,g)\mapsto v.g:=T\rho_{g}(v)$ (4.1)
and consider its restriction
$T_{e}G\times G\to TG,\quad(v,g)\mapsto v.g:=T\rho_{g}(v).$ (4.2)
Unlike on Lie groups, on half-Lie groups the latter action may not be smooth.
@@
would imply the smoothness of the multiplication map on $G$. In fact, since the point evaluation map $\varepsilon_{a}:AC_{L^{p}}([a,b],G)\to G$, $\eta\mapsto\eta(a)$ and the map $\zeta:G\to AC_{L^{p}}([a,b],G)$, $g\mapsto[t\mapsto g]$, are smooth (see [37]), the multiplication map on $G$ would be smooth as it coincides with the composition
@@
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space $E$ which admits local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],G)$. Let $\{t_{0},..,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ charts of $G$ such that verify the definition of absolute continuity for $\eta$. For each $i\in\{1,...,n\}$ we denote
$\eta_{i}:=\varphi_{i}\circ\eta_{[t_{i-1},t_{i}]}\in AC_{L^{p}}([t_{i-1},t_{i}],E).$
Then $\eta_{i}^{\prime}\in L^{p}([t_{i-1},t_{i}],E)$. Let $\eta^{\prime}_{i}=[\gamma_{i}]$ with $\gamma_{i}\in\mathcal{L}^{p}([t_{i-1},t_{i}],E)$. We define $\gamma:[a,b]\to TG$ via
$\gamma(t):=T(\varphi_{i})^{-1}(\eta_{i}(t),\gamma_{i}(t)),\text{ if }t\in[t_{i-1},t_{i}[\text{ with }i\in\{1,...,n\},$
and $\gamma(b)=T(\varphi_{n})^{-1}(\eta_{n}(b),\gamma_{n}(b))$. Then $\gamma$ is measurable (see [14]) and we define
$\dot{\eta}:=[\gamma].$
@@
${\dot{y}}(t)$ $=\gamma(t).y(t),\quad t\in[0,1]$ (4.3)
$y(0)$ $=e$ (4.4)
@@
$\operatorname{Evol}:L^{p}([0,1],T_{e}G)\to AC_{L^{p}}([a,b],G),\quad[\gamma]\mapsto\operatorname{Evol}(\gamma):=\eta_{\gamma}.$ (4.5)
@@
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space and $p\in[1,\infty]$. We say that $G$ is locally $L^{p}$-semiregular if there is exists an open 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that for each $[\gamma]\in B$ there exists a $AC_{L^{p}}$-Carathéodory solution $\eta_{\gamma}\in AC_{L^{p}}([0,1],G)$ of the equation
$\dot{y}(t)$ $=\gamma(t).y(t),\quad t\in[0,1]$
$y(0)$ $=e$
@@
For our purpose, we will use the subdivision property [33, Lemma 2.17].
@@
$\sup_{k\in\{0,...,n-1\}}q(\gamma_{n,k})\to 0,\quad\text{as }n\to\infty.$
@@
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. Then $G$ is locally $L^{p}$-semiregular if and only if $G$ is $L^{p}$-semiregular.
@@
Let $G$ be locally $L^{p}$-semiregular. Then there exists an 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that $\operatorname{Evol}(\beta)\in AC_{L^{p}}([a,b],G)$ exists for each $[\beta]\in B$. Without loss of generality we assume that
@@
For $[\beta]\in W_{\gamma}$ and $j\in\{0,...,n-1\}$ we write $\eta_{j}=\operatorname{Evol}(\beta_{n,j})$ and we define the function $\eta_{\beta}:[0,1]\to G$ via
@@
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. We say that $G$ is locally $L^{p}$-regular if $G$ is $L^{p}$-semiregular and there is exists a 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that its restricted evolution map $\operatorname{Evol}|_{B}$ is smooth.
The following lemma is just an application of [14, Proposition 5.25] or [32, Proposition 4.11] to our case.
@@
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space which admit a local addition and $p\in[1,\infty]$. We consider the evolution map with continuous values
@@
Then, $\operatorname{Evol}_{C}$ is continuous if and only if there exists a 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that the restricted evolution map $\operatorname{Evol}_{C}|_{B}$ is continuous.
@@
is continuous. Following Lemma 4.9, if $[\gamma]\in L^{p}([0,1],T_{e}G)$, for each $[\beta]\in W_{\gamma}$, we have that the construction $\operatorname{Evol}_{C}(\beta)$ implies that $\left(\text{Evol}_{C}\right)|_{W_{\gamma}}$ is just the product of composition of continuous maps. ∎
@@
Let $1\leq p<\infty$, $E_{1}$ be a Fréchet space, $E_{2}$ and $F$ be locally convex spaces and $U\subseteq E_{2}$ an open subset. If $r\in\mathbb{N}\cup\{0,\infty\}$ and $f:E_{1}\times U\to F$ is a $C^{r}$-map such that for each $y\in U$ the map $f(\cdot,y):E_{1}\to F$ is linear, then the map
@@
Hence $\tilde{f}(\gamma_{0},\eta_{0})\in L^{p}([0,1],F)$.
Since the space $C([0,1],E_{1})$ is dense in $L^{p}([0,1],E_{1})$ (see [14, Lemma A.4]) there exists a $\gamma_{c}\in C([0,1],E_{1})$ such that
$\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},\kappa}\leq\varepsilon/5.$
@@
With this, for each $\eta\in W$ and $\gamma\in L^{p}([0,1],E_{1})$ with $\lVert\gamma-\gamma_{0}\lVert_{L^{p},\kappa}\leq\varepsilon/5$ we estimate
$\left\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{0},\eta_{0})\right\lVert_{L^{p},\beta}$ $\leq\left\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{c},\eta)\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{0},\eta_{0})\right\lVert_{L^{p},\beta}$
$=\left\lVert\tilde{f}(\gamma-\gamma_{c},\eta)\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c},\eta_{0})-\tilde{f}(\gamma_{0},\eta_{0})\right\lVert_{L^{p},\beta}$
$\leq\left\lVert\tilde{f}(\gamma-\gamma_{c},\eta)\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c}-\gamma_{0},\eta_{0})\right\lVert_{L^{p},\beta}$
$\leq\Big{\lVert}\gamma-\gamma_{c}\Big{\lVert}_{L^{p},\kappa}+\left\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\right\lVert_{L^{p},\beta}+\Big{\lVert}\gamma_{c}-\gamma_{0}\Big{\lVert}_{L^{p},\kappa}.$
Since $\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},\kappa}\leq\varepsilon/5$ and $\lVert\gamma-\gamma_{0}\lVert_{L^{p},\kappa}\leq\varepsilon/5$ we have
$\lVert\gamma-\gamma_{c}\lVert_{L^{p},\kappa}\leq 2\varepsilon/5.$
Thus
$\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}\leq 2\varepsilon/5+2\varepsilon/5+\varepsilon/5=\varepsilon.$
Thus $\tilde{f}$ is continuous. By linearity in the first variable, the map $\overline{f}$ has a continuous differential in the first variable
$d_{1}\tilde{f}:L^{p}([0,1],E_{1})\times L^{p}([0,1],E_{1})\times C([0,1],E_{2})\to L^{p}([0,1],F).$
Let consider the map $f$ as a $C^{1}$-map. If $x\in E_{1}$, $y,y_{1}\in E_{2}$ and $t\in\mathbb{R}^{\times}$, then we have
@@
Hence the continuity of $\widetilde{d_{2}f}$ implies that the map $\widetilde{f}$ is $C^{1}$. Proceeding by induction, if $f$ is a $C^{k}$-map, since $\tilde{f}$ is linear in the first variable we have
@@
For the case $L_{rc}^{\infty}$ we recall [14, Proposition 2.3].
@@
Let $G$ be a right half-Lie group modeled in a Fréchet space $E$ which admits a local addition and $1\leq p<\infty$. Let $G$ be $L^{p}$-semiregular with continuous evolution map
$\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}G),\to C([0,1],G)\quad\gamma\mapsto\operatorname{Evol}_{C}(\gamma).$
@@
is continuous. If $G$ is a right half-Lie group modeled in an integral complete locally convex space $E$, then if we replace $L^{p}$ with $L_{rc}^{\infty}$ the result remains valid.
@@
Then each $f_{i}$ is continuous and for each $x\in V_{i}$ the function
$f_{i}(\cdot,x):T_{e}G\to E,\quad v\mapsto f_{i}(v,x)$
is linear. Hence by Lemma 4.12 the map
@@
$F:\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i})\to\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],E),\quad\left([\beta_{i}],\eta_{i}\right)_{i=1}^{n}\mapsto\left(f_{i}\circ([\beta_{i}],\eta_{i})\right)_{i=1}^{n}.$
@@
Let denote map